Assume the following for a firm that will operate for only three years. Year Zero: Capital...

  1. Assume the following for a firm that will operate for only three years. Year Zero: Capital Investment = $4M and increase in NWC = $0.2M; Years 1-3: NI = $3M, Depreciation =

$0.1M, Interest expense = $0.1M; Year 3: Recovery of earlier increase in NWC (i.e., decrease NWC by $0.2M), and in addition Cash Flow from Salvage = $3M. Suppose the tax rate is 30% for the firm. What is the IRR based on the FCFs for this firm in years 0 through 3?

  1. 79.03%
  2. 35.85%
  3. 71.26%
  4. 41.84%
Answer #1

C2 X for year o с year 0 year 1 $3,000,000 $ 100,000 $ 70,000 year 2 $3,000,000 $100,000 $70,000 year 3 $3,000,000 $100,000 $

for formulas and calculations, refer to the image below -

X for year o A year 0 year 1 3000000 100000 =100000*0.7 year 2 3000000 100000 =100000*0.7 year 3 3000000 100000 =100000*0.7 n

In case you have any query, kindly ask in comments.